In running the analysis for a multiple linear regression, you have two models with different number of variables (1st model with 3 variables and 2nd model with 4 variables), having 30 and 35 observations and R2 = 0.58 and 0.62, respectively.

Conduct an analysis to identify which model to be selected.
Answer #1

as we know that adjusted R2 =1-(1-R2)*(n-1)/(n-k-1)

therefore for first model:

adjusted R2 =1-(1-0.58)*(30-1)/(30-3-1) =0.5315

for second model:

adjusted R2 =1-(1-0.62)*(35-1)/(35-4-1) =0.5693

as adjusted R2 is higher for 2nd model therefore second model should be selected